OpenStudy (anonymous):

What is the first step to show that two lines are parallel, by looking at the equations of the lines, if they are written in standard form?

OpenStudy (anonymous):

Help pls

OpenStudy (anonymous):

They have same slope but different y-intercept. Look at the slope and y-intercept

OpenStudy (tkhunny):

You do need the slope. It's not obvious in "Standard Form". Ax + By = C ==> y = -(A/B)x + (C/B) There's your slope to compare. Find -A/B for each. 3x + 2y = 4 ==> Slope = -3/2 6x + 4y = 5 ==> Slope = -6/4 = -3/2 3x + y = 4 ==> Slope = -3 The top two examples COULD be parallel. If the constant is the same, that's a bit sneaky because they are really the SAME line. That's not really parallel.

OpenStudy (anonymous):

If both equations have the same slope they are parallel to each other. If one slope is a negative reciprocal of another, they are perpendicular.

OpenStudy (anonymous):

lines are parallel if they have the same slopes, but different y intercepts. So I would think that the first thing you would need to do is put your equations in y = mx + b form, where m is the slope. Once you find the slopes and y intercepts of the lines, compare them.